OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -8 and 512, respectively. The answers are: a_n = 2 * (-4)^n+1 a_n = 2 * 4^n-1 a_n = 2 * (-4)^n-1 a_n = 2 * 4^n I picked the first one for the answer. Is that right?

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen @mathmath333 help? I just need someone to check it.

OpenStudy (anonymous):

@mathstudent55 any help?

OpenStudy (anonymous):

@ant501 @gibbs @iambatman @ilovemathbutneedhelp @mathgeek27 @Shadowgirl15 HELP??!!

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

think it's A @sunshine201

OpenStudy (anonymous):

Thank you. Could you check one more? @ant501

OpenStudy (anonymous):

sure

OpenStudy (gibbs):

A

OpenStudy (anonymous):

open study acting up one sec

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6 The anwers are: a_n = -1856 + 98(n-1) a_n = -1856 - 98(n-1) a_n = -1856 - 98(n+1) a_n = -1856 + 98(n + 1) I picked A originally but I'm stuck between A and D. Which is it?

OpenStudy (anonymous):

I believe it's D

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

No that was all I had trouble with. Thank you! <3

OpenStudy (anonymous):

your welcome and thanx for the medal

OpenStudy (anonymous):

@chosenmatt